OpenStudy (anonymous):

find dy/dx for y=arctan(x/2)

OpenStudy (anonymous):

\[y=\tan^{-1} \frac{ x }{ a }\] \[\frac{ dy }{dx }=\frac{ 1 }{ 1+\frac{ x ^{2} }{a ^{2} } }*\frac{ 1 }{a }\]

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

yw